OpenStudy (nautika27):

A city is building a fence around a rectangular playground. If the perimeter of the playground is 90 feet and the length of one side of the fence is 30 feet, what is the width of the playground? A. 15 feet B. 45 feet C. 60 feet D. 30 feet

OpenStudy (nautika27):

@jhonyy9

OpenStudy (across):

(90-60)/2

OpenStudy (nautika27):

is / a divide sign ?

jhonyy9 (jhonyy9):

yes

OpenStudy (nautika27):

if so i got 60 @across

OpenStudy (nautika27):

oh ok

jhonyy9 (jhonyy9):

than 90 is the perimeter and you know that P=2l+2w and you know that l=30 so 90 = 2*30 +2w 90-60 =2w 30=2w w=30/2 = ?

OpenStudy (nautika27):

15

jhonyy9 (jhonyy9):

yes right

OpenStudy (nautika27):

thank you

jhonyy9 (jhonyy9):

np was my pleasure but do you understand it now ?

OpenStudy (nautika27):

yes :) and that will be all thanks for helping me

jhonyy9 (jhonyy9):

was my pleasure and anytime than i m here online just tag me there bye bye good luck